Shandy Corporation has a target capital structure of 40% debt and 60% common equity, with no preferred stock. Its before-tax cost of debt is 13%, and its marginal tax rate is 25%. The current stock price is P0 = $25.50. The last dividend was D0 = $2.50, and it is expected to grow at a 7% constant rate. What is its cost of common equity and its WACC? Do not round intermediate calculations. Round your answers to two decimal places.

rs = ___%

WACC =___ %
Answer #1

Answer :

Current price (P0 ) = D0 ( 1 + g ) / Ke - g

$25.50 = $2.50 ( 1 + 0.07 ) / Ke - 0.07

Ke - 0.07 = $2.675 / $ 25.50

Ke - 0.07 = 0.105

Ke = 0.105 + 0.07

Ke = 0.175

WACC =Ke * We + Kd ( 1 - T ) * Wd

Where We, Wd are weights in Equity & Debt respectively

Ke & Kd are Cost of equity & Cost of debt respectively

Post tax cost of debt = Cost of debt * ( 1 - T )

= 13% * ( 1 - 0.25 )

= 13% * 0.75

= 9.75%

WACC :

Source Weight Cost Weighted Cost
Debt 0.40 9.75% 3.9%
Equity 0.60 17.5% 10.5%
WACC => 14.4%

WACC = 14.4%

Cost of equity = 17.5%.